0 < 21. Round your 6. [10 points) Find all angles 0 such that 5coso - 2 = 2sin?and 0 answers to two decimal places.
Answer #1

Osoczn 5 cos o -2 =2 sinto Put sin?o=1-cos²o 560s 0 - 2=2(1-COSO) 2cos o + 5630 -4 = 0 compare to x = cosa - 2x2 + 5x -4 = 0......